A certain type of flashlight battery lasts on the average of 5 years with a standard deviation of 0.75 year. Assuming that the battery lives are normally distributed, find the probability that a given battery will last less than 3.5 years Find the probability


1
Expert's answer
2022-04-08T10:29:03-0400

"P(X<3.5)=P(Z<\frac{3.5-5}{0.75})=P(Z<-2)=0.0228."